OpenStudy (anonymous):

sinx+siny=1/3,answer siny-cos2x

OpenStudy (jamesj):

Is that sin(y) - cos(2x) or sin(y) - cos^2(x) ?

OpenStudy (anonymous):

yes

OpenStudy (jamesj):

Which one. The first one, or the second one. "Yes" is not an answer!

OpenStudy (anonymous):

i am sorry ,it is the first one

OpenStudy (anonymous):

can you answer this question

OpenStudy (jamesj):

Ok, now \[\cos (2x) = \cos^2 x - \sin^2 x = 1 - 2 \sin^2 x\] So\[\sin y - \cos(2x) = \sin y -1 + 2 \sin^2 x\]

OpenStudy (jamesj):

Now you can express sin x in terms of sin y. Substitute that expression into the last equation above and you'll have a quadratic equation in sin y

OpenStudy (anonymous):

assume that sin(y)-cos(2x)=z then sin(y) =z+cos(2x)=z+1-2sin^2(x) substitute in the first expression sin(x)+z+1-2sin^2(x)=1/3 sin^2(x)-.5sin(x)-(z+1)/2)=-1/6 (sin(x)-.25)^2-(z+1)/2=1/3 her (sin(x)-.25)^2 is in range L= [0:1.562] so L=(z+1)/2+1/3 L=z/2+5/6 z=2L-5/3 so z is in range [-1.6666:1.45833] z is not a number it is a set of number depends on the value of x

OpenStudy (jamesj):

That's wrong.